# Speedups in nonequilibrium thermal relaxation: Mpemba and related effects

### v1

@@ 4.2 Example: Granular smooth hard spheres
The simplest case of a dilute granular gas is a system composed of $N$ smooth, hard spheres of mass $m$ and diameter $\sigma$. We can analyze it within the framework presented above. Here, “hard” means that the particles can interact only through collisions; and when they collide, momentum is conserved, while part of the kinetic energy is lost according to the restitution coefficient $10<\alpha<$ 1. By “smooth,” we mean that there is no friction between the spheres during the collision. The angular velocity of each particle is then conserved and does not affect the linear velocities. This feature is accounted by the angular restitution coefficient $\beta=-1$. In this simple scenario, where the angular velocities play no role, it is less convenient to use the dimensionless parameters introduced in the previous section. We thus introduce a single relevant granular temperature $T$, which corresponds to $T_{t}$ in the notation of Sec. 4.1. In addition, moments that in general require three indices here need only a single index. Thus, $a_{i,j}^{(k)}$ is written as $a_{i}$, where $i$ refers to the order of the linear velocity. Nevertheless, the results outlined in the previous section still apply to the smooth, hard-sphere scenario.
@@
As outlined in the previous section, a stationary, finite-temperature state is maintained because particles are subject to an independent white-noise force with variance $m^{2}\chi_{0}^{2}$. (We restrict ourselves to homogeneous situations.) Therefore, the velocity distribution function reduces to $f(\vec{v},t)$ and obeys a Boltzmann-Fokker-Planck kinetic equation [151].
The time evolution of the granular temperature $T$ is coupled to that of the excess kurtosis (also referred to as the second Sonine coefficient). Specifically, $a_{2}$ which is the first Sonine correction in Eq. (89) to the Maxwellian distribution, simplifies dramatically to $a_{2}=\frac{3}{5}\langle v^{4}\rangle/\langle v^{2}\rangle^{2}-1$. From the kinetic equation for the moments of velocity distribution function, Eq. (96), one finds
@@
Figure 22: (a) Example of Mpemba effect and its inverse in a granular gas of smooth, hard spheres and Homogeneous Cooling State (b) Results from two simulation methods, the direct simulation Monte Carlo (DSMC) (open symbols) and molecular dynamics (MD) (filled symbols) data show excellent agreement with the theoretical prediction (lines); see (7.3). (c) Phase diagram of $\alpha$ and initial conditions. Figure reproduced from [4].
@@
The coupled equations for the temperature $\tilde{T}$ and $a_{2}$ can produce several types of nonmonotonic relaxation phenomena, including the direct, the inverse Mpemba effects and others [4, 153, 139, 136, 142]. By choosing the appropriate initial conditions for $a_{2}$ (initial state kurtosis), one can speed up or slow down the transient relaxation of a granular gas temperature for a given value of $T$. This is demonstrated in Fig. 22c, which shows the occurrence of the Mpemba effect as a function of the inelasticity parameter $\alpha$ and on the initial kurtosis and temperature. Note that, in this example, the effect is small in the range of possible choices of initial temperatures. Also, the initial temperatures in Fig. 22a have a maximum difference of about $0.01\%$. Of particular interest is Fig. 22a, where a non-thermostated granular gas is considered. This “Homogeneous Cooling State,” relaxes indefinitely towards the zero-temperature state, with $\chi_{0}=0$ continues to cool. In the homogeneous cooling state, the Mpemba effect is even smaller than in the heated case, but nevertheless, it exists; see inset Fig. 22. Note that, in contrast to other cases, the homogeneous cooling state has no natural scale for temperature. In this system, one thus uses an arbitrary reference temperature $T_{\mathrm{ref}}$. The rescaled variables are $T^{*}=T/T_{\mathrm{ref}}$ and $t^{*}=\kappa\sqrt{T_{\mathrm{ref}}}\,t$.
@@
Equation (103) shares strong similarities with the Markov framework discussed in Sec. 2, and one can indeed identify the existence of anomalous relaxations by analyzing the eigenvalues and eigenvectors of the corresponding matrix. Nevertheless, there are profound differences between the nature of the Mpemba effect in granular gases and in a general Markovian system. One such difference is that temperature in Eq. (103) is simultaneously one of the evolving quantities and what determines the existence of anomalous effect, whereas in the Markovian framework the existence of the effect is encoded in the projection of the initial condition onto the second-slowest left eigenvector of the rate matrix (see, e.g., Eq. 29). An additional distinction is that in granular gases the initial conditions are set by choosing both $T_{0}$ and $a_{2,0}$, so that both evolve in time towards a steady state: This follows from the fact that in granular gases the experimentally controlled parameters are $\alpha$ and $\chi$ – not $T$ and $a_{2}$. It is, therefore, a natural question to ask whether it is possible, in smooth granular gases, to tune the external parameter ($\chi$ and $\alpha$) such that the temperature is quenched, but $a_{2}$ does not change, and nevertheless to observe anomalous relaxation. Figure 22c directly answers this question: it shows that the quotient $\Delta\theta_{0}/\Delta a_{2,0}$ must remain inside the dashed area in order to observe the effect in the system discussed above. Therefore, in the specific model discussed here, when $a_{2,0}$ is kept fixed and only the steady-state temperature varies, the effect cannot appear (this can be seen in Fig. 22a for $\Delta a_{2,0}=0$). Note that the same happens in the viscoelastic model Sec. 4.4; see also Fig. 24b. This, however, is not a general property of granular gases. As discussed below, it can also happen in other models.
@@ 4.3 Example: Granular breaking equipartition
In a granular gas of rough spheres, the difference in initial temperatures that allows us to observe the Mpemba effect can be much larger [137]. As we discuss below, the mechanism that makes it larger is the coupling between the translational and rotational temperatures, which are of the same order of magnitude. This is related to the fact that granular gases do not reach an equilibrium state but rather a nonequilibrium steady state. As a consequence, rough granular gases do not follow the equilibrium equipartition theorem, and the energy in each degree of freedom depends on the values of the restitution coefficients, $\alpha$ and $\beta$. In addition, the basic features of the effect can be understood within the Maxwellian (Gaussian) approximation without resorting to higher-order cumulants. A similar reasoning was considered in [138] for rough disks.
@@
As explained more in detail in Sec. 4.1, in the long-time limit, the granular gas reaches a steady state that balances the action of the stochastic force against energy loss in collisions. For non-smooth granular gases, in which angular velocity plays an important role, this steady state is characterized by the temperature ratio $\theta^{\rm st}$ and total granular temperature $T^{\text{st}}$ in the Maxwellian approximation; see Eq. (82). In terms of the coefficients of restitution and the stochastic forcing intensity, the temperature ratio and the total granular temperature are (see [149]),
@@
where $\gamma^{\text{st}}$ is the dimensionless measure of noise intensity at stationarity; see Eq. (95). Note that $\theta^{\text{st}}\leq 1$ is independent of $\alpha$ in the Maxwellian approximation. Higher-order approximations introduce a weak dependence on $\alpha$, see [137].
@@
where we introduced the the constant $K\equiv\iota(1+\beta)^{2}/(1+\iota)^{2}$ and the auxiliary function
$\Phi(T^{*},\theta)=$ $\frac{2}{3}\left(\frac{\gamma^{\text{st}}}{T^{*}(1+\theta^{\text{st}})}-\sqrt{\frac{T^{*}(1+\theta^{\text{st}})}{\left(1+\theta\right)^{3}}}\gamma^{\text{st}}+K\sqrt{{\frac{T^{*}(1+\theta^{\text{st}})}{\left(1+\theta\right)^{3}}}}\left(1-\frac{\theta}{\theta^{\text{st}}}\right)(1-\theta^{\text{st}})\right)$ (109a)
accounting for the rate of change of (the logarithm of) the relative temperature. We underline here that the time evolution of $T^{*}$ is governed by a function that depends on both $T^{*}$ and $\theta$: as described below, this is a prerequisite for the emergence of a Mpemba effect; otherwise, no crossing between the temperature relaxation curves is possible.
To understand the mechanisms enabling a Mpemba effect, we consider two initial conditions $(T_{0A}^{*},\theta_{0A})$ and $(T_{0B}^{*},\theta_{0B})$, with $T_{0A}^{*}>T_{0B}^{*}>1$, for the same granular gas, i.e., with the same values of the coefficients of restitution $\alpha$, $\beta$, initial distribution function and the same stochastic force $\chi$. Let us denote by $T_{A}^{*}(\tau)$ and $T_{B}^{*}(\tau)$ the associated time evolution of the temperature to the steady state: A Mpemba effect exists in the system if there is a crossing time $\tau_{\times}$ such that $T_{A}^{*}(\tau)<T_{B}^{*}(\tau)$ for $\tau>\tau_{\times}$. Note that, in this framework, we do not demand that $\theta_{0A}=\theta_{0B}$, which again differs from the Markovian framework. In this system, it is not possible to have a crossing for $\theta_{0A}=\theta_{0B}$; see Fig. 23, right part of panel b.
@@
The most favorable conditions for the emergence of the Mpemba effect is when the kinetic energy is concentrated in (i) the translational velocity for the higher temperature $T^{*}_{0A}$, i.e., $\theta_{0A}=0$, and (ii) the rotational velocity for the lower temperature $T^{*}_{0B}$, i.e., $\theta_{0B}\rightarrow\infty$. Figure 23 shows the time evolution of the granular temperature, for $\alpha=0.7$ and $\alpha=0.9$, and $\beta=-0.5$, 00 and $\beta=0.5$. The Mpemba effect is clearly observed as the initially hottest sample (A) is the fastest one to reach the steady state.
@@
Figure 23: Example of large Mpemba effect in a granular gas of rough hard spheres, uniformly heated. (a) Time occurrence for two given values of $\beta$ and $\alpha$. Results are given by two simulation methods, the direct simulation Monte Carlo (DSMC) (open symbols) and molecular dynamics (MD) (filled symbols) data show an excellent agreement with the theoretical prediction (lines) see (7.3). (b) Phase diagram for the appearance of the effect for three different choices of initial conditions $(T^{*}_{0A},T^{*}_{0B})$ ($(i):(5,4)$, $(ii):(4,3)$, $(iii):(5,3)$), for each case below the lines the effect happens. (c) Shows the plane ($T^{*}_{0B}$,$T^{*}_{0A}/T^{*}_{0B}$). Below the plotted line the Mpemba effect may appear, provided that the kinetic energy of the initially hotter (colder) sample is sufficiently concentrated in the translational (rotational) degrees of freedom. The figures are reproduced from [137].
@@ 4.4 Example: Other granular models exhibiting anomalous relaxations
Viscoelastic granular particles – A more realistic model for granular gases is the viscoelastic granular particles model, for which the Mpemba effect was considered in [142]. As in the smooth-hard spheres case (Sec. 4.2), the spheres are smooth and angular velocities play no role. The main difference between the two models is the restitution coefficient, $\alpha$. For smooth, hard spheres, it is assumed to be constant, whereas for viscoelastic granular particles, it depends on velocity. The velocity dependence is modeled by $\alpha=\alpha(v)\backsimeq 1-\gamma_{v}|\vec{v}_{1,2}\cdot\vec{\sigma}_{1,2}|^{1/5}+(3/5)\gamma_{v}^{2}|\vec{v}_{1,2}\cdot\vec{\sigma}_{1,2}|^{3/5}$ [154]. Note that, in this case, the difference in initial temperatures that allows us to observe the Mpemba effect can be slightly larger than for smooth, hard spheres; see Fig. 24a. Also, the second cumulant ($a_{3}=\frac{4}{5}\langle v^{4}\rangle/\langle v^{2}\rangle^{2}-\frac{8}{105}\langle v^{6}\rangle/\langle v^{2}\rangle^{3}-2$) is significant and must be taken into account. Interestingly, the Kovacs effect, which we discuss in Sec. 9.1, was also observed in this specific model.
@@
Figure 24: a) Example of Mpemba and its inverse in a granular gas of viscoelastic spheres. Note that the difference in initial temperatures is slightly larger than in the hard sphere case. Results are given by two simulation methods, the direct simulation Monte Carlo (DSMC) (dotted lines with crosses) and molecular dynamics (MD) (open symbols). The data show an excellent agreement with the theoretical predictions (lines); see (7.3). b) Phase diagram as a function of $\gamma_{v}$ and initial conditions. These figure are adapted from [142].
@@
Maxwell model – A different simplification that allows to study smooth hard spheres systems is the Maxwell model. In this approximation, the collision frequency is not a function of the particle velocity, but rather is a fixed parameter of the system. It affects significantly the collision kernel. A series of manuscripts considered the Mpemba and related effects in this and related models, [71, 155, 156] as well as driven monodispersed and bidispersed granular gases [61]. Interestingly, the Mpemba effect exists in monodispersed gas only if both temperature and additional parameters are quenched; however, in bidispersed granular systems, the Mpemba effect exists even if only the temperature is quenched, as in the Markovian framework. Moreover, the latter case contains a strong version of the Mpemba effect: for specially chosen initial temperatures, the asymptotic relaxation has a different exponential rate [61].
@@
Figure 25: (a) Sketch of the two ME protocols used in an inertial suspension. The protocol for FQE (From quasi equilibrium intital state, dashed lines) and FS (from a sheared steady initial state, solid line). For the FQE the shear rate and the environmental temperature are changed at $t=0$, (left) from 00 to $\dot{\gamma}_{tar}$, (right) from $T_{ini}$ environmental to $T_{tar}$ environmental respectively. For FS, (left) the shear rate is changed at t = 0 from $\dot{\gamma}_{ini}$ to $\dot{\gamma}_{tar}$ while (right) the environmental temperature is kept at $T_{tar}$ environmental. (b) Left: Diagram of the evolutions of the temperatures from the equilibrium (top) and from the nonequilibrium (bottom). Right: illustrates the temperature differences during the (top) normal Mpemba effect (NME), (middle) anomalous Mpemba effect (AME), and (bottom) NME + AME. (c) Numerical simulations show the two types of effects, where $\theta_{i}=T/T_{env}$ is the dimensionless temperature and $i=FS,FQE$ indicates the protocol. The left panel shows the $\theta_{i}$ evolution, and the right panel shows the difference of $\theta_{FQE}-\theta_{FS}$ where anomalous and normal Mpemba effects can be observed. The figures are adapted from [141].
@@
Inertial suspension of granular particles – Another granular system where the normal and what the authors denote as “anomalous Mpemba effect” (as well as its inverse) were found is an inertial suspension of granular particles under shear [141]. The system that was considered is a suspension in a shear flow, namely in the presence of a fluid flow with velocity field given by $\vec{u}=(\dot{\gamma}y,\vec{u}_{\perp}=0)$, and subject to a uniform kinetic temperature $T$. Each suspension particle is therefore described by the following Langevin equation
@@
where $\vec{p}_{i}=m(\vec{v}_{i}-\dot{\gamma}y_{i}\vec{e}_{x})$ with $\vec{e}_{x}$ is a unit vector in the $x$ direction, $\vec{F}_{i}^{imp}$ expresses the impulsive force coming from collisions between grains, and $\vec{\xi}_{i}(t)=\xi_{i,q}\vec{e}_{q}$ is delta correlated noise from the surrounding thermal fluid: $\langle\vec{\xi}_{i}\rangle=0$ and $\langle\xi_{i,q}(t)\xi_{j,s}(t^{\prime})\rangle=2\zeta T_{\rm env}\delta_{i\!j}\delta_{qs}\delta(t,-t^{\prime})$, where indexes $i,j$ refer to particle numbers and $q,s={x,y,z}$. This model, therefore, has two external parameters: the environmental temperature $T_{\rm env}$, and the shear rate $\dot{\gamma}$. These quantities together determine the (nonequilibrium) granular temperature. Within this framework, the authors considered “normal” and “anomalous” Mpemba effects. For simplicity, we use these terms in this specific subsection.
The “normal” Mpemba effect stems from the following result, that can be shown from the above equation: when a high enough temperature equilibrium initial condition (here equilibrium is the steady state corresponding to no shearing, $\dot{\gamma}=0$) is compared to a nonequilibrium initial condition (the steady state of a system with $\dot{\gamma}\neq 0$) with the same initial temperature, then one finds that the nonequilibrium initial condition relaxes faster, namely $\dot{T}_{\rm eq}(t=0)<\dot{T}_{\rm neq}(t=0)$. This implies that for small-enough difference in initial temperature, the nonequilibrium initial condition can cool faster and overcome the equilibrium initial condition. Note that this effect concerns the initial stage of the relaxation, in contrast to the Markovian effect that concerns the long-time limit.

### v3

@@ 4.2 Example: Granular smooth hard spheres
The simplest case of a dilute granular gas is a system composed of $N$ smooth, hard spheres of mass $m$ and diameter $\sigma$. We can analyze it within the framework presented above theoretically and numerically by means of molecular dynamics (MD) and direct simulation Monte Carlo (DSMC) methods (see Sec.  7.3). Here, “hard” means that the particles can interact only through collisions; and when they collide, momentum is conserved, while part of the kinetic energy is lost, as quantified by the restitution coefficient $10<\alpha<$ 1. By “smooth,” we mean that there is no friction between the spheres during the collision. The angular velocity of each particle is then conserved and does not affect the linear velocities. This feature is accounted by the angular restitution coefficient $\beta=-1$. In this simple scenario, where the angular velocities play no role, it is less convenient to use the dimensionless parameters introduced in the previous section. We thus introduce a single relevant granular temperature $T$, which corresponds to $T_{t}$ in the notation of Sec. 4.1. In addition, moments that in general require three indices here need only a single index. Thus, $a_{i,j}^{(k)}$ is written as $a_{i}$, where $i$ refers to the order of the linear velocity. Nevertheless, the results outlined in the previous section still apply to the smooth, hard-sphere scenario.
@@
As outlined in the previous section, a stationary, finite-temperature state is maintained because particles are subject to an independent white-noise force with variance $m^{2}\chi_{0}^{2}$. (We restrict ourselves to homogeneous situations.) Therefore, the velocity distribution function reduces to $f(\vec{v},t)$ and obeys a Boltzmann-Fokker-Planck kinetic equation [158, 159].
The time evolution of the granular temperature $T$ is coupled to that of the excess kurtosis, also referred to as the second Sonine coefficient. Specifically, $a_{2}$ which is the first Sonine correction in Eq. (89) to the Maxwellian distribution, simplifies dramatically to $a_{2}=\frac{3}{5}\langle v^{4}\rangle/\langle v^{2}\rangle^{2}-1$. From the kinetic equation for the moments of velocity distribution function, Eq. (96), one finds
@@
Figure 22: (a) Example of Mpemba effect and its inverse in a granular gas of smooth, hard spheres and homogeneous cooling state (b) Results from two simulation methods, DSMC (open symbols) and MD (filled symbols) data show excellent agreement with the theoretical prediction (lines); see (7.3). (c) Phase diagram of $\alpha$ and initial conditions. Source: Reprinted with permission from [4].
@@
The coupled equations for the temperature $\tilde{T}$ and $a_{2}$ can produce several types of nonmonotonic relaxation phenomena, including the direct, the inverse Mpemba effects and others [4, 161, 146, 143, 149]. By choosing the appropriate initial conditions for $a_{2}$ (initial state kurtosis), one can speed up or slow down the transient relaxation of a granular gas temperature for a given value of $T$. This is demonstrated in Fig. 22c, which shows the occurrence of the Mpemba effect as a function of the inelasticity parameter $\alpha$ and on the initial kurtosis and temperature. Note that, in this example, the effect is small in the range of possible choices of initial temperatures. Also, the initial temperatures in Fig. 22a have a maximum difference of about $0.01\%$. Of particular interest is Fig. 22 b), where a non-thermostated granular gas is considered. This “homogeneous cooling state” relaxes indefinitely towards the zero-temperature state, with $\chi_{0}=0$ continues to cool. In the homogeneous cooling state, the Mpemba effect is even smaller than in the heated case, but nevertheless, it exists; see Fig. 22 b). Note that, in contrast to other cases, the homogeneous cooling state has no natural scale for temperature. In this system, one thus uses an arbitrary reference temperature $T_{\mathrm{ref}}$. The rescaled variables are $T^{*}=T/T_{\mathrm{ref}}$ and $t^{*}=\kappa\sqrt{T_{\mathrm{ref}}}\,t$.
@@
Equation (103) shares strong similarities with the Markov framework discussed in Sec. 2, and one can indeed identify the existence of anomalous relaxations by analyzing the eigenvalues and eigenvectors of the corresponding matrix. Nevertheless, there are profound differences between the nature of the Mpemba effect in granular gases and in a general Markovian system. One such difference is that temperature in Eq. (103) is both one of the evolving quantities and also determines the existence of anomalous effect. By contrast, in the Markovian framework, the existence of the effect is encoded in the projection of the initial condition onto the second-slowest left eigenvector of the rate matrix (see, e.g., Eq. (28)). An additional distinction is that in granular gases the initial conditions are set by choosing both $T_{0}$ and $a_{2,0}$, so that both evolve in time towards a steady state: This follows from the fact that in granular gases the experimentally controlled parameters are $\alpha$ and $\chi$, not $T$ and $a_{2}$. For smooth granular gases, it is therefore natural to ask whether one can tune the external parameters ($\chi$ and $\alpha$) to quench the temperature while leaving $a_{2}$ unchanged, and nevertheless to observe anomalous relaxation. Figure 22 c) directly answers this question: it shows that the quotient $\Delta\theta_{0}/\Delta a_{2,0}$ must remain inside the dashed area in order to observe the effect in the system discussed above. Therefore, in the specific model discussed here, when $a_{2,0}$ is kept fixed and only the steady-state temperature varies, the effect cannot appear (this can be seen in Fig. 22 a) for $\Delta a_{2,0}=0$). Note that the same happens in the viscoelastic model Sec. 4.4; see also Fig. 24 b). This, however, is not a general property of granular gases. As discussed below, it can also happen in other models.
@@ 4.3 Example: Granular breaking equipartition
In a granular gas of rough spheres, the difference in initial temperatures that allows us to observe the Mpemba effect can be much larger [144]. As we discuss below, the mechanism that makes it larger is the coupling between the translational and rotational temperatures, which are of the same order of magnitude. This is related to the fact that granular gases do not reach an equilibrium state but rather a nonequilibrium steady state. Rough granular gases then do not follow the equilibrium equipartition theorem, and the energy in each degree of freedom depends on the values of the restitution coefficients, $\alpha$ and $\beta$. In addition, the basic features of the effect can be understood within the Maxwellian (Gaussian) approximation without resorting to higher-order cumulants. A similar reasoning was considered in [145] for rough disks.
@@
As explained more in detail in Sec. 4.1, in the long-time limit, the granular gas reaches a steady state that balances the action of the stochastic force against energy loss in collisions. For non-smooth granular gases, in which angular velocity plays an important role, this steady state is characterized by the temperature ratio $\theta^{\rm st}$ and total granular temperature $T^{\text{st}}$ in the Maxwellian approximation; see Eq. (82). In terms of the coefficients of restitution and the stochastic forcing intensity, the temperature ratio and the total granular temperature are, see [156],
@@
where $\gamma^{\text{st}}$ is the dimensionless measure of noise intensity at stationarity; see Eq. (95). Note that $\theta^{\text{st}}\leq 1$ is independent of $\alpha$ in the Maxwellian approximation. Higher-order approximations introduce a weak dependence on $\alpha$; see [144].
@@
where we introduced the constant $K\equiv\iota(1+\beta)^{2}/(1+\iota)^{2}$ and the auxiliary function
$\Phi(T^{*},\theta)=\frac{2}{3}\left(\frac{\gamma^{\text{st}}}{T^{*}(1+\theta^{\text{st}})}-\sqrt{\frac{T^{*}(1+\theta^{\text{st}})}{\left(1+\theta\right)^{3}}}\gamma^{\text{st}}+K\sqrt{{\frac{T^{*}(1+\theta^{\text{st}})}{\left(1+\theta\right)^{3}}}}\left(1-\frac{\theta}{\theta^{\text{st}}}\right)(1-\theta^{\text{st}})\right),$ (109)
accounting for the rate of change of (the logarithm of) the relative temperature. We underline here that the time evolution of $T^{*}$ is governed by a function that depends on both $T^{*}$ and $\theta$. As described below, this is a prerequisite for the emergence of a Mpemba effect; otherwise, no crossing between the temperature relaxation curves is possible.
To understand the mechanisms enabling a Mpemba effect, we consider two initial conditions $(T_{0A}^{*},\theta_{0A})$ and $(T_{0B}^{*},\theta_{0B})$, with $T_{0A}^{*}>T_{0B}^{*}>1$, for the same granular gas, i.e., with the same values of the coefficients of restitution $\alpha$, $\beta$ and the same stochastic force $\chi$. Let us denote by $T_{A}^{*}(\tau)$ and $T_{B}^{*}(\tau)$ the associated time evolution of the temperature to the steady state. A Mpemba effect then exists in the system if there is a crossing time $\tau_{\times}$ such that $T_{A}^{*}(\tau)<T_{B}^{*}(\tau)$ for $\tau>\tau_{\times}$. Note that, in this framework, we do not demand that $\theta_{0A}=\theta_{0B}$, which again differs from the Markovian framework. In this system, it is not possible to have a crossing for $\theta_{0A}=\theta_{0B}$; see Fig. 23, right part of panel b.
@@
The most favorable conditions for the emergence of the Mpemba effect is when the kinetic energy is concentrated in (i) the translational velocity for the higher temperature $T^{*}_{0A}$, i.e., $\theta_{0A}=0$, and (ii) the rotational velocity for the lower temperature $T^{*}_{0B}$, i.e., $\theta_{0B}\rightarrow\infty$. Figure 23 shows the time evolution of the granular temperature, for $\alpha=0.7$ and $\beta=-0.5$. The Mpemba effect is clearly observed as the initially hottest sample (a) is the fastest one to reach the steady state.
@@
Figure 23: Example of large Mpemba effect in a granular gas of rough hard spheres, uniformly heated. (a) Time occurrence for two given values of $\beta$ and $\alpha$. Results are given by two simulation methods, DSMC (open symbols) and MD (filled symbols) data show an excellent agreement with the theoretical prediction (lines) see (7.3). (b) Phase diagram for the appearance of the effect for three different choices of initial conditions $(T^{*}_{0A},T^{*}_{0B})$ ($(i)$: $(5,4)$, $(ii)$: $(4,3)$, $(iii)$: $(5,3)$), for each case below the lines the effect happens. (c) Shows the plane ($T^{*}_{0B}$,$T^{*}_{0A}/T^{*}_{0B}$). Below the plotted line the Mpemba effect may appear, provided that the kinetic energy of the initially hotter (colder) sample is sufficiently concentrated in the translational (rotational) degrees of freedom. Source: Reprinted with permission from [144].
@@ 4.4 Example: Other granular models exhibiting anomalous relaxations
Viscoelastic granular particles. A more realistic model for granular gases is the viscoelastic granular particles model, for which the Mpemba effect was considered in [149]. As in the smooth-hard spheres case (Sec. 4.2), the spheres are smooth and angular velocities play no role. The main difference between the two models is the restitution coefficient, $\alpha$. For smooth, hard spheres, it is assumed to be constant, whereas for viscoelastic granular particles, it depends on velocity. The velocity dependence is modeled by $\alpha=\alpha(v)\backsimeq 1-\gamma_{v}|\vec{v}_{1,2}\cdot\vec{\sigma}_{1,2}|^{1/5}+(3/5)\gamma_{v}^{2}|\vec{v}_{1,2}\cdot\vec{\sigma}_{1,2}|^{3/5}$ [162]. Note that, in this case, the difference in initial temperatures that allows us to observe the Mpemba effect can be slightly larger than for smooth, hard spheres; see Fig. 24a. Also, the third cumulant ($a_{3}=\frac{4}{5}\langle v^{4}\rangle/\langle v^{2}\rangle^{2}-\frac{8}{105}\langle v^{6}\rangle/\langle v^{2}\rangle^{3}-2$) is significant and must be taken into account. Interestingly, the Kovacs effect, which we discuss in Sec. 9.1, was also observed in this specific model.
@@
Figure 24: (a) Example of Mpemba and its inverse in a granular gas of viscoelastic spheres. Note that the difference in initial temperatures is slightly larger than in the hard sphere case. Results are given by two simulation methods, DSMC (dotted lines with crosses) and MD (open symbols). The data show an excellent agreement with the theoretical predictions (lines); see (7.3). (b) Phase diagram as a function of $\gamma_{v}$ and initial conditions. Source: Adapted with permission from [149].
@@
Maxwell model. A different simplification that allows to study smooth hard spheres systems is the Maxwell model. In this approximation, the collision frequency is not a function of the particle velocity, but rather is a fixed parameter of the system. It affects significantly the collision kernel. A series of manuscripts considered the Mpemba and related effects in this and related models, [163, 164, 165], as well as driven monodispersed and bidispersed granular gases [61]. Interestingly, the Mpemba effect exists in monodispersed gas only if both temperature and additional parameters are quenched; however, in bidispersed granular systems, the Mpemba effect exists even if only the temperature is quenched, as in the Markovian framework. Moreover, the latter case contains a strong version of the Mpemba effect: for specially chosen initial temperatures, the asymptotic relaxation has a different exponential rate [61].
@@
Figure 25: (a) Sketch of the two ME protocols used in an inertial suspension. The protocol for FQE (from a quasi-equilibrium initial state, dashed lines) and FS (from a sheared steady initial state, solid line). For the FQE, the shear rate and the environmental temperature are changed at $t=0$, (left) from 00 to $\dot{\gamma}_{tar}$, (right) from $T_{ini}$ environmental to $T_{tar}$ environmental respectively. For FS, (left) the shear rate is changed at t = 0 from $\dot{\gamma}_{ini}$ to $\dot{\gamma}_{tar}$ while (right) the environmental temperature is kept at $T_{tar}$ environmental. (b) Left: Diagram of the evolutions of the temperatures from the equilibrium (top) and from the nonequilibrium (bottom). Right: illustrates the temperature differences during the (top) normal Mpemba effect (NME), (middle) anomalous Mpemba effect (AME), and (bottom) NME + AME. (c) Numerical simulations show the two types of effects, where $\theta_{i}=T/T_{\rm env}$ is the dimensionless temperature and $i=$ FS,FQE indicates the protocol. The left panel shows the $\theta_{i}$ evolution, and the right panel shows the difference $\theta_{\rm FQE}-\theta_{\rm FS}$, where anomalous and normal Mpemba effects can be observed. Source: Adapted with permission from [148].
@@
Inertial suspension of granular particles. Another granular system where the normal and what the authors denote as “anomalous Mpemba effect” (as well as its inverse) were found is an inertial suspension of granular particles under shear [148]. The system that was considered is a suspension in a shear flow, namely in the presence of a fluid flow with velocity field given by $\vec{u}=(\dot{\gamma}y,\vec{u}_{\perp}=0)$, and subject to a uniform kinetic temperature $T$. Each suspension particle is therefore described by the Langevin equation
@@
where $\vec{p}_{i}=m(\vec{v}_{i}-\dot{\gamma}y_{i}\vec{e}_{x})$ with $\vec{e}_{x}$ is a unit vector in the $x$ direction, $\vec{F}_{i}^{imp}$ expresses the impulsive force coming from collisions between grains, and $\vec{\xi}_{i}(t)=\xi_{i,q}\vec{e}_{q}$ is delta correlated noise from the surrounding thermal fluid: $\langle\vec{\xi}_{i}\rangle=0$ and $\langle\xi_{i,q}(t)\xi_{j,s}(t^{\prime})\rangle=2\zeta T_{\rm env}\delta_{i\!j}\delta_{qs}\delta(t,-t^{\prime})$, where indices $i,j$ refer to particle numbers and $q,s={x,y,z}$. This model, therefore, has two external parameters, the environmental temperature $T_{\rm env}$, and the shear rate $\dot{\gamma}$. These quantities together determine the (nonequilibrium) granular temperature. Note that the Enskog equation [154] is used here at finite density to describe the system and the influence of the packing fraction on the ME is also analyzed. Within this framework, the authors considered “normal” and “anomalous” Mpemba effects. For simplicity, we use these terms in this specific subsection.
The “normal” Mpemba effect stems from a result which can be shown from the above equation: when a high-enough temperature equilibrium initial condition (here equilibrium is the steady state corresponding to no shearing, $\dot{\gamma}=0$) is compared to a nonequilibrium initial condition (the steady state of a system with $\dot{\gamma}\neq 0$) with the same initial temperature, then the nonequilibrium initial condition relaxes faster, namely $\dot{T}_{\rm eq}(t=0)<\dot{T}_{\rm neq}(t=0)$. This implies that for small-enough difference in initial temperature, the nonequilibrium initial condition can cool faster and overcome the equilibrium initial condition. Note that this effect concerns the initial stage of the relaxation, in contrast to the Markovian effect that concerns the long-time limit.